# CRAG - Comprehensive RAG Benchmark

### v1

@@ A.4.1 Quality
We send the CRAG public test set question as input to each of the SOTA RAG systems and collect the responses for human grading. Note that the original Query Time and the provided retrieval results in CRAG are not used in this setting. We simply test the questions and ask human graders to grade the responses based on when the query was made to the SOTA system. We called Copilot Pro, Gemini Advanced, and ChatGPT Plus through their web interfaces and Perplexity.ai through its API. We set the conversation style to “Precise” when calling Copilot Pro and the temperature to 00 when calling Perplexity.ai. We select GPT-4o and llama-3-sonar-large-32k-online as the base LLM when calling ChatGPT Plus and Perplexity.ai, respectively.
@@ A.4.2 Latency
As the latency measured by API call and web interface interactions are not directly comparable, we further called Perplexity.ai through its web interface. The latency under this setting is 4,629ms. However, the latency number measured here may not correspond to the accuracy numbers from the API calls.

### v2

@@ A.4.1 Quality
We send the CRAG public test set question as input to each of the SOTA RAG systems and collect the responses for human grading. Note that the original Query Time and the provided retrieval results in CRAG are not used in this setting. We simply test the questions and ask human graders to grade the responses based on when the query was made to the SOTA system. We called Copilot Pro, Gemini Advanced, and ChatGPT Plus through their web interfaces and Perplexity.ai through its API. Meta SG, designed as a smart glasses (SG) assistant, includes default on-device components such as Automatic Speech Recognition (ASR) and Text-to-Speech (TTS), which are not typically enabled by default in other systems. To ensure a fair comparison, we excluded these on-device components, ensuring that answer quality was not affected. We called each system on the following dates in Pacific Time: 05/12/2024$\sim$05/16/2024 (Copilot Pro), 05/20/2024$\sim$05/28/2024 (Gemini Advanced), 05/27/2024$\sim$06/02/2024 (ChatGPT Plus), 05/15/2024$\sim$05/16/2024 (Perplexity.ai), and 07/02/2024 (Meta SG). We set the conversation style to “Precise” when calling Copilot Pro and the temperature to 00 when calling Perplexity.ai. We select GPT-4o and llama-3-sonar-large-32k-online as the base LLM when calling ChatGPT Plus and Perplexity.ai, respectively.
@@ A.4.2 Latency
The latency of Perplexity.ai measured via API call is 2,455ms. Since latency measured by API call and web interface interactions are not directly comparable, we further called Perplexity.ai through its web interface and reported the latency under this setting in Table 6. Note that this latency may not correspond to the accuracy numbers from the API calls. For Meta SG, we estimated a latency comparable to other web interface interactions by excluding on-device components such as ASR and TTS from the overall end-to-end latency measurement.
A.5 Limitations
The three tasks in our benchmark do not directly evaluate the construction of a first-stage retrieval candidate pool, a demanding retrieval task in its own right. This design decision ensures that the competition remains both challenging and achievable within the KDD Cup’s required three-month timeframe. Despite the limitation, users of our dataset have the option to use the union of all 220K web pages as a corpus to build a retriever. While this corpus does not match the entire web, it allows for fair comparisons and manageable costs.